A trolley,mass 5kg moves at 4m.s east across a frictionless horizontal surface. A brick of mass 1.5kg is dropped onto the trolley . Calculate the change in momentum of the 5kg trolley

1
Expert's answer
2022-04-19T09:59:56-0400

"mv=(m+m_1)u,"

"u=\frac{mv}{m+m_1}=3.1~\frac ms."